# Orbifold modifications of complex analytic varieties

### v4

@@ 7. Modification around an irreducible singular locus
Let $X$ be a complex analytic variety with klt singularities. Assume that $S\subseteq X_{\operatorname{\textsubscript{\rm sing}}}$ is a compact irreducible component of codimension 2 in $X$. Then there is a projective bimeromorphic morphism $f\colon Y\to X$, with $Y$ normal, such that the following properties hold. Let $\Delta$ be the whole reduced exceptional divisor of $f$.
@@
    The divisorial critical locus of $\psi_{i}\colon W_{i}\to W$ is contained in $\Delta$.
@@
We consider the data $X$, $\mathfrak{X}$, $S$ and $H$ as follows. Let $X$ be a complex analytic variety. Assume that $S\subseteq X_{\operatorname{\textsubscript{\rm sing}}}$ is a compact irreducible component of codimension 2. We are only interested in the behavior of some neighborhood of $S$ in $X$. Let $H\subseteq S$ be a proper closed subset, and let $\mathfrak{X}=(X_{i},G_{i},\pi_{i})_{i\in I}$ be a complex analytic orbispace, with a finite family $I$ of orbifold charts, whose quotient space is $X$. We set $S_{i}=\pi_{i}^{-1}(S)$ and $H_{i}=\pi_{i}^{-1}(H)$. Assume that $S_{i}$ is an irreducible component of the singular locus of $X_{i}$, and that $X_{i}$ has the same type of canonical singularities at points of $S_{i}\setminus H_{i}$. Furthermore, the intersection of the divisorial critical locus of $X_{i}\to X$ and $S$ is contained in $H$.
We will prove Proposition 7.1 in several steps, and the sketch is as follows. Since $X$ has klt singularities, there is a complex analytic orbispace $(X_{i},G_{i})$ induced by the local index-one covers of $\omega_{X}$, see Example 2.12. The difficult case is when $X_{i}$ is singular around the preimage of $S$. To study this case, we may assume the situation of Setup 7.2. We first apply Proposition 4.1 on each orbifold chart $X_{i}$. Then we can obtain some projective bimeromorphic morphism $f\colon Y\to X$. We then locate on a neighborhood $U$ of the strict transform $S_{Y}$ of $S$ in $Y$. Replacing $X$ by $U$ and $S$ by $S_{Y}$, we can assume that each $X_{i}$ satisfies the conclusion of Proposition 4.1. As a result, we are able to apply Proposition 5.2 on each $X_{i}$. Afterward, we may assume that each $X_{i}$ satisfies the conclusion of Proposition 5.2. Now we can apply Proposition 6.7 on each $X_{i}$. This will complete the proof of Proposition 7.1.
@@
With the notation of Setup 7.2, there is a projective bimeromorphic morphism $f\colon Y\to X$ such that the following properties hold.
@@
For each $i$, we let $f_{i}\colon Y_{i}\to X_{i}$ be the construction of Proposition 4.1, with respect to $H_{i}\subseteq S_{i}\subseteq X_{i}$. Since the construction of Proposition 4.1 is functorial, we see that there is a natural action of $G_{i}$ on $Y_{i}$, and the family $(Y_{i},G_{i})_{i\in I}$ defines a complex analytic orbispace $\mathfrak{Y}$. Let $Y$ be its quotient space. Then there is an induced bimeromorphic morphism $f\colon Y\to X$. Since each step in the construction of $f_{i}$ is the blowup of some ideal sheaf, which is compatible along the overlaps, there is a $f_{i}$-exceptional and $f_{i}$-ample divisor $E_{i}$ which is compatible along the overlaps. The collection $(E_{i})_{i\in I}$ induces a $f$-exceptional and $f$-ample divisor in $Y$. This proves the items (1), (2) and (4).
@@
It remains to prove the item (5). To this end, we can apply Lemma 2.16 to the orbi-divisor $((H_{Y})_{i})_{i\in I}$ in the complex orbifold $((S_{Y})_{i},G_{i})_{i\in I}$. Since $(H_{Y})_{i}$ is a snc divisor in $(S_{Y})_{i}$, the construction of Lemma 2.16 is to successively blowing up certain strata of $(H_{Y})_{i}$, which are compatible along the overlaps. Hence we can blow up $Y_{i}$ at the same centers, see Subsection 3.B. These blowups are again compatible along the overlaps. Let $Y^{\prime}_{i}\to Y_{i}$ be this blowup. Then we obtain a complex analytic orbi-space $(Y_{i}^{\prime},G_{i})_{i\in I}$ with quotient space $Y^{\prime}$. Replacing $Y$ by $Y^{\prime}$, we can assume that $H_{Y}\subseteq S_{Y}$ induces a snc orbi-divisor on the orbifold $((S_{Y})_{i},G_{i})_{i\in I}$. This proves the item (5) and completes the proof of the lemma. ∎
@@
With the notation of Setup 7.2, we assume the properties of Lemma 7.4 hold. Then there is a projective bimeromorphic morphism $f\colon Y\to X$ such that the following properties hold.
@@
    The indeterminacy locus of $f^{-1}$ is a proper closed subset of $S$. In particular, we can define the strict transform $S_{Y}$ of $S$ in $Y$, and the preiamge $H_{Y}$ of $H$ in $S_{Y}$ under the morphism $f|_{S_{Y}}\colon S_{Y}\to S$.
@@
    The divisorial critical locus of $Y_{i}\to Y$ is contained in $\Delta_{Y}$.
@@
where $X_{i,j+1}\to X_{i,j}$ is the blowup of $X_{i,j}$ at $D^{j+1}_{i}$ for $N$ times, as in Subsection 3.C. Then there is a natural action of $G_{i}$ on $Y_{i}$. Moreover, the family $(Z_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace $\mathfrak{Z}$ with quotient space $Y$. If $\psi_{i}\colon Z_{i}\to X_{i}$ is the composition of the sequence above, then the collection $(\psi_{i})_{i\in I}$ induces a proper bimeromorphic morphism $f\colon Y\to X$. In addition, since $\psi_{i}$ is the composition of blowups whose centers are compatible along the overlaps, we deduce that there is some $\psi_{i}$-ample and $\psi_{i}$-exceptional divisor which is compatible along the overlaps. Hence there is a $f$-exceptional and $f$-ample divisor. Let $(S_{Z})_{i}$ be the strict transform of $S_{i}$ in $Z_{i}$. By construction, there is some open neighborhood $U_{i}$ of $(S_{Z})_{i}$, such that $U_{i}$ is isomorphic to an open neighborhood of $\{\mathbf{0}_{3}\}\times(S_{Z})_{i}$ in the hypersurface in $\mathbb{\mathbb{D}}^{3}\times(S_{Z})_{i}$ defined by a function as in Proposition 5.2. Since the intersection of the divisorial critical locus of $X_{i}\to X$ and $S$ is contained in $H$, up to enlarging $N$, we can assume that the divisorial critical locus of $Z_{i}\to Y$ is contained in $\Delta_{Y}$. We set $\widehat{G}_{i}=G_{i}$ and $\widehat{V}_{i}=X_{i}$ in this case.
[…]
@@
Now, as in the Case 1, we blow up $\widehat{V_{i}}$ at the preimage of each irreducible component of $H_{W}$ for $N$ times as in Subsection 3.C. We obtain a complex analytic orbispace $\mathfrak{Z}=(Z_{i},\widehat{G}_{i})$ with quotient space $Y^{\prime}$. Let $(S_{Z})_{i}$ be the strict transform of $(S_{\widehat{V}})_{i}$ in $Z_{i}$. By the construction of Lemma 5.9, there is some open neighborhood $U_{i}$ of $(S_{Z})_{i}$ in $Z_{i}$, such that $U_{i}$ is isomorphic to an open neighborhood of $\{\mathbf{0}_{3}\}\times(S_{Z})_{i}$ in the hypersurface in $\mathbb{\mathbb{D}}^{3}\times(S_{Z})_{i}$ defined by a function as in Proposition 5.2 with respect to the pair $((S_{Z})_{i},(H_{Z})_{i})$. Furthermore, there is a proper bimeromorphic morphism $h\colon Y^{\prime}\to\widehat{V}$ with some $h$-ample and $h$-exceptional divisor. Since the indeterminacy locus of $h^{-1}$ is contained in the compact set $S_{W}$, we deduce that $h$ extends to a projective bimeromorphic morphism $h\colon Y\to W$ by Lemma 2.3. Hence we obtain a projective bimeromorphic morphism $f\colon Y\to X$, with a $f$-exceptional and $f$-ample divisor. Furthermore, up to enlarging $N$, we can assume that the divisorial critical locus of $Z_{i}\to Y$ is contained in $\Delta_{Y}$.
[…]
@@
The divisorial critical locus of $Y_{i}\to V^{\prime}_{i}$ is contained in the $\psi_{i}$-exceptional locus, hence is contained in the preimages of $\Delta_{Y}$ in $V^{\prime}_{i}$. Since the divisorial critical locus of $V^{\prime}_{i}\to Y$ is contained in $\Delta_{Y}$, we deduce the item (6). This completes the proof of the lemma. ∎
@@
With the notation of Setup 7.2, let $\Delta$ be a reduced divisor in $X$ and let $\Delta_{i}$ be its preimage in $X_{i}$. Suppose that $\Delta$ contains the divisorial critical locus of $X_{i}\to X$, and that $H$ induces a snc orbi-divisor on the complex orbifold $(S_{i},G_{i})_{i\in I}$. Assume that the data $X_{i}$, $\Delta_{i}$, $S_{i}$, $H_{i}$ and $G_{i}$ satisfy the items (1) and (2) of Setup 6.6. Then, there is a projective bimeromorphic morphism $f\colon Y\to X$ with $Y$ normal such that the following properties hold.
@@
    The indeterminacy locus of $f^{-1}$ is a proper closed subset of $S$. In particular, we can define the strict transform $S_{Y}$ of $S$ in $Y$, and the preiamge $H_{Y}$ of $H$ in $S_{Y}$ under the morphism $f|_{S_{Y}}\colon S_{Y}\to S$.
@@
    The divisorial critical locus of $\psi_{k}\colon\overline{W}_{k}\to W$ is contained in $\Delta_{Y}$, where $\Delta_{Y}=f^{-1}(\Delta)$.
@@
More concretely, for each point $x\in S_{i}$, we let $S_{x}$ be an open neighborhood of it as the one in the item (4) of Setup 6.6. In particular, if $(G_{i})_{x}\subseteq G_{i}$ is the subgroup which fixes $x$, then $(S_{x},(G_{i})_{x})$ is an orbifold chart of $\pi_{i}(S_{x})$, compatible with the orbifold $(S_{i},G_{i})_{i\in I}$. There is a positive integer $N_{x}$, such that Proposition 6.7 holds on $X_{i}$ with respect to $S_{x}$ and $N_{x}$. Since $S$ is compact, by considering $S_{x}$ for all $x\in S_{i}$ and all $i\in I$, we can obtain a finite family $(x_{k})_{k\in K}$ of points and an application $\sigma\colon K\to I$ such that the following properties hold. Each point $x_{k}$ belongs to $S_{\sigma(k)}$ and
@@
Let $D^{1},...,D^{e}$ be the irreducible components of $H$. Since $H$ induces a snc orbi-divisor on the complex analytic orbispace $(S_{i},G_{i})_{i\in I}$, if the preimage $D^{j}_{i}$ of $D^{j}$ in $X_{i}$ is not empty, then it is smooth and hence irreducible. Let $f_{i}\colon Y_{i}\to X_{i}$ be the morphism obtained by blowing up of $X_{i}$ at $D^{j}_{i}$ for $6N$ times, successively for $j=1,...,e$, as in Subsection 3.C. Then there is a natural action of $G_{i}$ on $Y_{i}$ and the collection $(Y_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace $\mathfrak{Y}$ with quotient space $Y$. Furthermore, there is an induced projective bimeromorphic morphism $f\colon Y\to X$ with a $f$-ample and $f$-exceptional divisor. Let $S_{Y}$ be the strict transform of $S$ in $Y$.
From the choice of $N$, for each $k\in K$, there is an open neighborhood $W_{k}$ of the strict transform of $S_{x_{k}}$ in $Y_{\sigma(k)}$, which is invariant under $G_{k}:=(G_{\sigma(k)})_{x_{k}}$, such that there is a finite cover $\nu_{k}\colon\overline{W}_{k}\to W_{k}$ satisfying the conclusions of Proposition 6.7. The property (4) of Proposition 6.7 implies that the natural morphism $\psi_{k}\colon\overline{W}_{k}\to Y$ is finite and Galois of group $\overline{G}_{k}$ over the normalization of its image. We remark that, up to shrinking $W_{k}$ around $S_{x_{k}}$, we can assume that $(W_{k},G_{k})$ is an orbifold chart over $Y$ compatible with $\mathfrak{Y}$.
Let $W$ be the union of the images of the $W_{k}$’s in $Y$. Then $W$ is the quotient space of the complex analytic orbispace $(W_{k},G_{k})_{k\in K}$ and is an open neighborhood of $S_{Y}$. By Lemma 4.8, up to shrinking $W_{k}$, we can assume that $W_{k}$ is locally irreducible. Hence $W$ is locally irreducible as well. Thus the normalization morphism $W^{\prime}\to W$ is a homeomorphism, and we may identify $W^{\prime}$ and $W$ as the same topological space. We will show that $(\overline{W}_{k},\overline{G}_{k})_{k\in K}$ induces a complex analytic orbispace, whose quotient space is $W^{\prime}$.
@@
Hence there is an open neighborhood $X^{\prime}$ of $f(w)$ in $X$, and an orbifold chart $(Z,G^{\prime})$ of $X^{\prime}$, such that there are inclusions of orbifold charts $\zeta_{k}\colon Z\to X_{\sigma(k)}$ and $\zeta_{l}\colon Z\to X_{\sigma(l)}$. Let $(Z_{Y},G^{\prime})$ be the orbifold chart on $f^{-1}(X^{\prime})$ induced by $(Z,G^{\prime})$, which is compatible with $\mathfrak{Y}$. Then there is an open neighborhood $U$ of $w$ in $\psi_{k}(\overline{W}_{k})\cap\psi_{l}(\overline{W}_{l})$, and an orbifold chart $(V,G)$ over $U$, such that there are inclusions of orbifold charts
$\iota_{k}\colon V\to W_{k},\ \iota_{l}\colon V\to W_{l}\mbox{ and }V\to Z_{Y}.$

### v6

@@ 7. Modification around an irreducible singular locus
Let $X$ be a complex analytic variety with klt singularities. Assume that $S\subseteq X_{\operatorname{\textsubscript{\rm sing}}}$ is a compact irreducible component of codimension 2 in $X$. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$, with $Y$ normal, such that the following properties hold. Let $\Delta$ be the whole reduced exceptional divisor of $f$.
@@
    The divisorial critical locus of $\psi_{i}\colon W_{i}\rightarrow W$ is contained in $\Delta$.
@@
We consider the data $X$, $\mathfrak{X}$, $S$ and $H$ as follows. Let $X$ be a complex analytic variety. Assume that $S\subseteq X_{\operatorname{\textsubscript{\rm sing}}}$ is a compact irreducible component of codimension 2. We are only interested in the behavior of some neighborhood of $S$ in $X$. Let $H\subseteq S$ be a proper closed subset, and let $\mathfrak{X}=(X_{i},G_{i},\pi_{i})_{i\in I}$ be a complex analytic orbispace, with a finite family $I$ of orbifold charts, whose quotient space is $X$. We set $S_{i}=\pi_{i}^{-1}(S)$ and $H_{i}=\pi_{i}^{-1}(H)$. Assume that $S_{i}$ is an irreducible component of the singular locus of $X_{i}$, and that $X_{i}$ has the same type of canonical singularities at points of $S_{i}\setminus H_{i}$. Furthermore, the intersection of the divisorial critical locus of $X_{i}\rightarrow X$ and $S$ is contained in $H$.
We will prove Proposition 7.1 in several steps, and the sketch is as follows. Since $X$ has klt singularities, there is a complex analytic orbispace $(X_{i},G_{i})$ induced by the local index-one covers of $\omega_{X}$, see Example 2.12. The difficult case is when $X_{i}$ is singular around the preimage of $S$. To study this case, we may assume the situation of Setup 7.2. We first apply Proposition 4.1 on each orbifold chart $X_{i}$. Then we can obtain some projective bimeromorphic morphism $f\colon Y\rightarrow X$. We then locate on a neighborhood $U$ of the strict transform $S_{Y}$ of $S$ in $Y$. Replacing $X$ by $U$ and $S$ by $S_{Y}$, we can assume that each $X_{i}$ satisfies the conclusion of Proposition 4.1. As a result, we are able to apply Proposition 5.2 on each $X_{i}$. Afterward, we may assume that each $X_{i}$ satisfies the conclusion of Proposition 5.2. Now we can apply Proposition 6.7 on each $X_{i}$. This will complete the proof of Proposition 7.1.
@@
With the notation of Setup 7.2, there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ such that the following properties hold.
@@
For each $i$, we let $f_{i}\colon Y_{i}\rightarrow X_{i}$ be the construction of Proposition 4.1, with respect to $H_{i}\subseteq S_{i}\subseteq X_{i}$. Since the construction of Proposition 4.1 is functorial, we see that there is a natural action of $G_{i}$ on $Y_{i}$, and the family $(Y_{i},G_{i})_{i\in I}$ defines a complex analytic orbispace $\mathfrak{Y}$. Let $Y$ be its quotient space. Then there is an induced bimeromorphic morphism $f\colon Y\rightarrow X$. Since each step in the construction of $f_{i}$ is the blowup of some ideal sheaf, which is compatible along the overlaps, there is a $f_{i}$-exceptional and $f_{i}$-ample divisor $E_{i}$ which is compatible along the overlaps. The collection $(E_{i})_{i\in I}$ induces a $f$-exceptional and $f$-ample divisor in $Y$. This proves the items (1), (2) and (4).
@@
It remains to prove the item (5). To this end, we can apply Lemma 2.16 to the orbi-divisor $((H_{Y})_{i})_{i\in I}$ in the complex orbifold $((S_{Y})_{i},G_{i})_{i\in I}$. Since $(H_{Y})_{i}$ is a snc divisor in $(S_{Y})_{i}$, the construction of Lemma 2.16 is to successively blowing up certain strata of $(H_{Y})_{i}$, which are compatible along the overlaps. Hence we can blow up $Y_{i}$ at the same centers, see Subsection 3.B. These blowups are again compatible along the overlaps. Let $Y^{\prime}_{i}\rightarrow Y_{i}$ be this blowup. Then we obtain a complex analytic orbi-space $(Y_{i}^{\prime},G_{i})_{i\in I}$ with quotient space $Y^{\prime}$. Replacing $Y$ by $Y^{\prime}$, we can assume that $H_{Y}\subseteq S_{Y}$ induces a snc orbi-divisor on the orbifold $((S_{Y})_{i},G_{i})_{i\in I}$. This proves the item (5) and completes the proof of the lemma. ∎
@@
With the notation of Setup 7.2, we assume the properties of Lemma 7.4 hold. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ such that the following properties hold.
@@
    The indeterminacy locus of $f^{-1}$ is a proper closed subset of $S$. In particular, we can define the strict transform $S_{Y}$ of $S$ in $Y$, and the preiamge $H_{Y}$ of $H$ in $S_{Y}$ under the morphism $f|_{S_{Y}}\colon S_{Y}\rightarrow S$.
@@
    The divisorial critical locus of $Y_{i}\rightarrow Y$ is contained in $\Delta_{Y}$.
@@
where $X_{i,j+1}\rightarrow X_{i,j}$ is the blowup of $X_{i,j}$ at $D^{j+1}_{i}$ for $N$ times, as in Subsection 3.C. Then there is a natural action of $G_{i}$ on $Y_{i}$. Moreover, the family $(Z_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace $\mathfrak{Z}$ with quotient space $Y$. If $\psi_{i}\colon Z_{i}\rightarrow X_{i}$ is the composition of the sequence above, then the collection $(\psi_{i})_{i\in I}$ induces a proper bimeromorphic morphism $f\colon Y\rightarrow X$. In addition, since $\psi_{i}$ is the composition of blowups whose centers are compatible along the overlaps, we deduce that there is some $\psi_{i}$-ample and $\psi_{i}$-exceptional divisor which is compatible along the overlaps. Hence there is a $f$-exceptional and $f$-ample divisor. Let $(S_{Z})_{i}$ be the strict transform of $S_{i}$ in $Z_{i}$. By construction, there is some open neighborhood $U_{i}$ of $(S_{Z})_{i}$, such that $U_{i}$ is isomorphic to an open neighborhood of $\{\mathbf{0}_{3}\}\times(S_{Z})_{i}$ in the hypersurface in $\mathbb{\mathbb{D}}^{3}\times(S_{Z})_{i}$ defined by a function as in Proposition 5.2. Since the intersection of the divisorial critical locus of $X_{i}\rightarrow X$ and $S$ is contained in $H$, up to enlarging $N$, we can assume that the divisorial critical locus of $Z_{i}\rightarrow Y$ is contained in $\Delta_{Y}$. We set $\widehat{G}_{i}=G_{i}$ and $\widehat{V}_{i}=X_{i}$ in this case.
Case 2. We assume that the singularities of $X_{i}$ at $S_{i}\setminus H_{i}$ are of $D$-type. We will proceed as in Lemma 5.9 on each $X_{i}$. Let $\varphi_{i}\colon W_{i}\rightarrow X_{i}$ be the composition of blowups at the centers $D^{1}_{i},...,D^{k}_{i}$ successively. Then there is a natural action of $G_{i}$ on $W_{i}$ and $(W_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace $\mathfrak{W}$ with quotient space $W$. Furthermore, there is an induced projective bimeromorphic morphism $\varphi\colon W\rightarrow X$ with a $\varphi$-ample and $\varphi$-exceptional divisor. Let $S_{W}$ be the strict transform of $S$ in $W$, and let $(S_{W})_{i}$ be the strict transforms of $S_{i}$ in $W_{i}$. Let $H_{W}$ be the preimage of $H$ in $S_{W}$, and let $(H_{W})_{i}$ the preimage of $H_{W}$ in $W_{i}$. Let $\mathcal{I}(D^{j}_{i})$ be the ideal sheaf of $D^{j}_{i}$ in $X_{i}$. There is an open neighborhood $V_{i}$ of $(S_{W})_{i}$ such that $\varphi_{i}^{-1}\mathcal{I}(D^{j}_{i})\cdot\mathcal{O}_{V_{i}}$ is generated by a holomorphic function $\sigma^{j}$ on $V_{i}$, for j=1,,kj=1,...,k, see Remark 3.5 and Remark 5.10. Shrinking $V_{i}$ we may assume that it is $G_{i}$-invariant. We note that for every j=1,,kj=1,...,k, the collection of subspaces $(D^{j}_{i})_{i\in I}$ is compatible along the overlaps. By applying Lemma 2.13 for $k$ times, up to shrinking $V_{i}$ and up to refining the family $I$, we can assume the following properties. Let $\widehat{V}_{i}=V_{i}[\sqrt[6]{\sigma_{1}},...,\sqrt[6]{\sigma_{k}}]$. Then the morphism $\widehat{V}_{i}\rightarrow Z$ is finite and Galois over its image. Let $\widehat{G}_{i}$ be its Galois group. The collection $(\widehat{V_{i}},\widehat{G_{i}})_{i\in I}$ induces a complex analytic orbispace $\widehat{\mathfrak{V}}$, whose quotient space $\widehat{V}$ is an open neighborhood of $S_{W}$ in $W$. Let $(S_{\widehat{V}})_{i}$ be the preimage of $(S_{W})_{i}$ in $\widehat{V}_{i}$. Then the finite morphism $(S_{\widehat{V}})_{i}\rightarrow(S_{W})_{i}$ is a cyclic cover, branched over the components of $(H_{W})_{i}$. Indeed, $(S_{\widehat{V}})_{i}$ is also isomorphic to a contractible neighborhood of the origin of $\mathbb{C}^{n}$, and the preimage $(H_{\widehat{V}})_{i}$ of $(H_{W})_{i}$ in $(S_{\widehat{V}})_{i}$ is isomorphic to the union of some coordinate hyperplanes. In particular, $H_{W}$ induces a snc orbi-divisor on the complex orbifold $((S_{\widehat{V}})_{i},\widehat{G}_{i})_{i\in I}$.
@@
Now, as in the Case 1, we blow up $\widehat{V_{i}}$ at the preimage of each irreducible component of $H_{W}$ for $N$ times as in Subsection 3.C. We obtain a complex analytic orbispace $\mathfrak{Z}=(Z_{i},\widehat{G}_{i})$ with quotient space $Y^{\prime}$. Let $(S_{Z})_{i}$ be the strict transform of $(S_{\widehat{V}})_{i}$ in $Z_{i}$. By the construction of Lemma 5.9, there is some open neighborhood $U_{i}$ of $(S_{Z})_{i}$ in $Z_{i}$, such that $U_{i}$ is isomorphic to an open neighborhood of $\{\mathbf{0}_{3}\}\times(S_{Z})_{i}$ in the hypersurface in $\mathbb{\mathbb{D}}^{3}\times(S_{Z})_{i}$ defined by a function as in Proposition 5.2 with respect to the pair $((S_{Z})_{i},(H_{Z})_{i})$. Furthermore, there is a proper bimeromorphic morphism $h\colon Y^{\prime}\rightarrow\widehat{V}$ with some $h$-ample and $h$-exceptional divisor. Since the indeterminacy locus of $h^{-1}$ is contained in the compact set $S_{W}$, we deduce that $h$ extends to a projective bimeromorphic morphism $h\colon Y\rightarrow W$ by Lemma 2.3. Hence we obtain a projective bimeromorphic morphism $f\colon Y\rightarrow X$, with a $f$-exceptional and $f$-ample divisor. Furthermore, up to enlarging $N$, we can assume that the divisorial critical locus of $Z_{i}\rightarrow Y$ is contained in $\Delta_{Y}$.
In both cases, we have constructed a projective bimeromorphic morphism $f\colon Y\rightarrow X$, and a complex analytic orbispace $\mathfrak{Z}=(Z_{i},\widehat{G}_{i})$, whose quotient space, denoted by $Z$, is an open neighborhood of the strict transform $S_{Y}$ of $S$ in $Y$. By construction, the items (1) and (2) of the lemma hold for $f$. Furthermore, the items (3) and (4) hold for $\mathfrak{Z}=(Z_{i},\widehat{G}_{i})$. However, in the item (5), there is an issue which is the item (2) of Setup 6.6. To handle it, we apply the method of cyclic covers as in Lemma 6.13. Since $\psi_{i}\colon Z_{i}\rightarrow\widehat{V}_{i}$ is obtained by blowing up the components $(D_{\widehat{V}})_{i}^{j}$ of $(H_{\widehat{V}})_{i}$, we can proceed as in the first paragraph of the Case 2 above. By Lemma 2.13, up to refining the family $I$, there is a $\widehat{G}_{i}$-invariant neighborhood $V^{\prime}_{i}$ of $(S_{Z})_{i}$ in $Z_{i}$, and a cyclic covering space $Y_{i}=V^{\prime}_{i}[\sqrt[6]{\delta_{1}},...,\sqrt[6]{\delta_{k}}]$, such that $\delta_{j}$ is a generator of $\psi_{i}^{-1}\mathcal{I}((D_{\widehat{V}})_{i}^{j})\cdot\mathcal{O}_{V^{\prime}_{i}}$, and that $Y_{i}\rightarrow V^{\prime}_{i}/\widehat{G}_{i}$ is Galois of group $\overline{G}_{i}$. Then the collection $(Y_{i},\overline{G}_{i})_{i\in I}$ defines a complex analytic orbispace $\mathfrak{Y}$, whose quotient space is an open neighborhood of $S_{Y}$. Moreover, by Lemma 6.13 and Remark 3.5, the item (4) holds on $Y_{i}$, and the vanishing orders of the functions $\alpha$ and $\beta$ in the defining function of $U_{i}$ in $\mathbb{D}^{3}\times(S_{Y})_{i}$ are divisible by 6 along each component of $(H_{Y})_{i}$. We notice that the item (1) of Setup 6.6 is satisfied by the construction of $U_{i}^{\prime}$. Hence we obtain the item (5) of the present lemma.
@@
The divisorial critical locus of $Y_{i}\rightarrow V^{\prime}_{i}$ is contained in the $\psi_{i}$-exceptional locus, hence is contained in the preimages of $\Delta_{Y}$ in $V^{\prime}_{i}$. Since the divisorial critical locus of $V^{\prime}_{i}\rightarrow Y$ is contained in $\Delta_{Y}$, we deduce the item (6). This completes the proof of the lemma. ∎
@@
With the notation of Setup 7.2, let $\Delta$ be a reduced divisor in $X$ and let $\Delta_{i}$ be its preimage in $X_{i}$. Suppose that $\Delta$ contains the divisorial critical locus of $X_{i}\rightarrow X$, and that $H$ induces a snc orbi-divisor on the complex orbifold $(S_{i},G_{i})_{i\in I}$. Assume that the data $X_{i}$, $\Delta_{i}$, $S_{i}$, $H_{i}$ and $G_{i}$ satisfy the items (1) and (2) of Setup 6.6. Then, there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ with $Y$ normal such that the following properties hold.
@@
    The indeterminacy locus of $f^{-1}$ is a proper closed subset of $S$. In particular, we can define the strict transform $S_{Y}$ of $S$ in $Y$, and the preiamge $H_{Y}$ of $H$ in $S_{Y}$ under the morphism $f|_{S_{Y}}\colon S_{Y}\rightarrow S$.
@@
    The divisorial critical locus of $\psi_{k}\colon\overline{W}_{k}\rightarrow W$ is contained in $\Delta_{Y}$, where $\Delta_{Y}=f^{-1}(\Delta)$.
@@
More concretely, for each point $x\in S_{i}$, we let $S_{x}$ be an open neighborhood of it as the one in the item (4) of Setup 6.6. In particular, if $(G_{i})_{x}\subseteq G_{i}$ is the subgroup which fixes $x$, then $(S_{x},(G_{i})_{x})$ is an orbifold chart of $\pi_{i}(S_{x})$, compatible with the orbifold $(S_{i},G_{i})_{i\in I}$. There is a positive integer $N_{x}$, such that Proposition 6.7 holds on $X_{i}$ with respect to $S_{x}$ and $N_{x}$. Since $S$ is compact, by considering $S_{x}$ for all $x\in S_{i}$ and all $i\in I$, we can obtain a finite family $(x_{k})_{k\in K}$ of points and an application $\sigma\colon K\rightarrow I$ such that the following properties hold. Each point $x_{k}$ belongs to $S_{\sigma(k)}$ and
@@
Let $D^{1},...,D^{e}$ be the irreducible components of $H$. Since $H$ induces a snc orbi-divisor on the complex analytic orbispace $(S_{i},G_{i})_{i\in I}$, if the preimage $D^{j}_{i}$ of $D^{j}$ in $X_{i}$ is not empty, then it is smooth and hence irreducible. Let $f_{i}\colon Y_{i}\rightarrow X_{i}$ be the morphism obtained by blowing up of $X_{i}$ at $D^{j}_{i}$ for $6N$ times, successively for j=1,,ej=1,...,e, as in Subsection 3.C. Then there is a natural action of $G_{i}$ on $Y_{i}$ and the collection $(Y_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace $\mathfrak{Y}$ with quotient space $Y$. Furthermore, there is an induced projective bimeromorphic morphism $f\colon Y\rightarrow X$ with a $f$-ample and $f$-exceptional divisor. Let $S_{Y}$ be the strict transform of $S$ in $Y$.
From the choice of $N$, for each $k\in K$, there is an open neighborhood $W_{k}$ of the strict transform of $S_{x_{k}}$ in $Y_{\sigma(k)}$, which is invariant under $G_{k}:=(G_{\sigma(k)})_{x_{k}}$, such that there is a finite cover $\nu_{k}\colon\overline{W}_{k}\rightarrow W_{k}$ satisfying the conclusions of Proposition 6.7. The property (4) of Proposition 6.7 implies that the natural morphism $\psi_{k}\colon\overline{W}_{k}\rightarrow Y$ is finite and Galois of group $\overline{G}_{k}$ over the normalization of its image. We remark that, up to shrinking $W_{k}$ around $S_{x_{k}}$, we can assume that $(W_{k},G_{k})$ is an orbifold chart over $Y$ compatible with $\mathfrak{Y}$.
Let $W$ be the union of the images of the $W_{k}$’s in $Y$. Then $W$ is the quotient space of the complex analytic orbispace $(W_{k},G_{k})_{k\in K}$ and is an open neighborhood of $S_{Y}$. By Lemma 4.8, up to shrinking $W_{k}$, we can assume that $W_{k}$ is locally irreducible. Hence $W$ is locally irreducible as well. Thus the normalization morphism $W^{\prime}\rightarrow W$ is a homeomorphism, and we may identify $W^{\prime}$ and $W$ as the same topological space. We will show that $(\overline{W}_{k},\overline{G}_{k})_{k\in K}$ induces a complex analytic orbispace, whose quotient space is $W^{\prime}$.
@@
Hence there is an open neighborhood $X^{\prime}$ of $f(w)$ in $X$, and an orbifold chart $(Z,G^{\prime})$ of $X^{\prime}$, such that there are inclusions of orbifold charts $\zeta_{k}\colon Z\rightarrow X_{\sigma(k)}$ and $\zeta_{l}\colon Z\rightarrow X_{\sigma(l)}$. Let $(Z_{Y},G^{\prime})$ be the orbifold chart on $f^{-1}(X^{\prime})$ induced by $(Z,G^{\prime})$, which is compatible with $\mathfrak{Y}$. Then there is an open neighborhood $U$ of $w$ in $\psi_{k}(\overline{W}_{k})\cap\psi_{l}(\overline{W}_{l})$, and an orbifold chart $(V,G)$ over $U$, such that there are inclusions of orbifold charts
$\iota_{k}\colon V\rightarrow W_{k},\ \iota_{l}\colon V\rightarrow W_{l}\mbox{ and }V\rightarrow Z_{Y}.$